OpenStudy (kirbykirby):

X, Y are iid random variables both with an exp(1) distribution. Find P(X/Y > t) for t>0

OpenStudy (zarkon):

\[P\left(\frac{X}{Y}>t\right)=P(X>tY)\] \[=\mathbb{E}_{Y}\left[P(X>tY|Y=y)\right]\] \[=\cdots\]